# Neural Common Neighbor with Completion for Link Prediction

### v3

@@ 1 Introduction
However, models combining SF and MPNN still have much room for improvement. They can generally be concluded into two architectures: SF-then-MPNN and SF-and-MPNN, as shown in Figure 2. SEAL (Zhang & Chen, 2018) adds target-link-specific hand-crafted features to the node features of the input graphs of MPNN, whose output node representations are then pooled to produce link representations. SEAL belongs to the SF-then-MPNN architecture, which leverages SF to augment the input graph of MPNN. Though SF-then-MPNN models achieve provably high expressivity (Zhang et al., 2021), they require running MPNN on a different graph for each target link, resulting in significant computational overhead.
@@
In contrast, Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) decouple the structural feature from MPNN. They directly incorporate manually created pairwise features with individual node representations produced by MPNN as link representations, necessitating only a single run of MPNN on the original graph. These models fall under the SF-and-MPNN category, where structural features and MPNN are independent. Such methods have limited expressivity. For example, SEAL can capture the representations of common neighbor nodes, while BUDDY can only count the number of common neighbors.
To solve the drawback of the two architectures above, we propose MPNN-then-SF architecture, which initially applies MPNN to the original graph and then uses structural features to guide the pooling of node representations. This approach offers strong expressivity and scalability: similar to SF-then-MPNN models, MPNN-then-SF can capture the node features of common neighbors, and similar to SF-and-MPNN models, it runs MPNN only once for all target links. We introduce the Neural Common Neighbor (NCN) as an instantiation of the MPNN-then-SF architecture. In experiments, NCN outperforms existing models in both scalability and performance.
@@
  • We introduce the Neural Common Neighbor (NCN) for link prediction using the MPNN-then-SF architecture, demonstrating superior performance and scalability compared to existing models.
@@ 2 Preliminary
We consider an undirected graph ${\mathcal{G}}=(V,E,A,X)$, where $V=\{1,2,\ldots,n\}$ represents a set of $n$ nodes, $E\subseteq V\times V$ denotes the set of edges, $X\in\mathbb{R}^{n\times F}$ is a node feature matrix whose $v$-th row $X_{v}$ is the feature of node $v$, and adjacency matrix $A\in\mathbb{R}^{n\times n}$ is a symmetric matrix, whose $(u,v)$ element is $1$ if $(u,v)\in E$ and 00 otherwise. The degree of node $u$ is $d(u,A):=\sum_{v=1}^{n}A_{uv}$. Node $u$’s neighbors are nodes connected to $u$, $N(u,A):=\{v|v\!\in\!V,A_{uv}>0\}$. For simplicity of notations, we use $N(u)$ to denote $N(u,A)$ when $A$ is fixed. Common neighbor means nodes connected to both $i$ and $j$: $N(i)\cap N(j)$.
@@
We define $A^{l}$ as a high-order adjacency matrix, where $A_{uv}^{l}$ represents the number of walks of length $l$ between nodes $u$ and $v$ in graph $A$. $N(u,A^{l})=\{v|v\in V,A^{l}_{uv}>0\}$ denotes the set of nodes connected to $u$ by a walk of length $l$ in graph $A$, equivalent to the neighbors in higher-order adjacency. $N_{l}(u,A)$ denotes the set of nodes whose shortest path distance to $u$ in graph $A$ is $l$. Existing works define high-order neighbors as either $N(u,A^{l})$ or $N_{l}(u,A)$. More generally, the neighborhood of $u$ can be expressed as $N_{l_{1}}(u,A^{l_{2}})$, returning all nodes with a shortest path distance of $l_{1}$ to $u$ in the high-order graph $A^{l_{2}}$. We use $N_{l_{1}}^{l_{2}}(u)$ to denote $N_{l_{1}}(u,A^{l_{2}})$ when $A$ is fixed. Given a target link $(i,j)$, their general neighborhood overlap is given by $N_{l_{1}}^{l_{2}}(i)\cap N_{l_{1}^{\prime}}^{l_{2}^{\prime}}(j)$, and their neighborhood difference is given by $N_{l_{1}}^{l_{2}}(i)\!-\!N_{l_{1}^{\prime}}^{l_{2}^{\prime}}(j)$.
@@ 3 Related Work
3.1 Link Prediction Model
There are three primary categories of link prediction models. Node embedding methods (Perozzi et al., 2014; Grover & Leskovec, 2016; Tang et al., 2015) map each node to an embedding vector and combine the embeddings of target nodes to predict link. Link prediction heuristics (Liben-Nowell & Kleinberg, 2003; Barabási & Albert, 1999; Zhou et al., 2009; Adamic & Adar, 2003) develop hand-crafted structural features. GNNs utilize Graph Neural Networks to predict link existence. Among these GNNs, Graph Autoencoder (GAE) (Kipf & Welling, 2016) uses the inner product of the MPNN representations of two target nodes, $\langle\text{MPNN}(i,A,X),\text{MPNN}(j,A,X)\rangle$, as the representations of link $(i,j)$. It uses MPNN only and thus fails to capture pairwise relations between nodes. In contrast, various GNNs combining MPNN and structural features (Zhang & Chen, 2018; Yun et al., 2021; Chamberlain et al., 2023) have achieved state-of-the-art performance. Take SEAL (Zhang & Chen, 2018) as an example. For a target link $(i,j)$, SEAL initially augments node feature $X$ by concatenating each node’s shortest path distance to $(i,j)$ and extracts a $k$-hop subgraph from the whole graph, generating augmented node feature $X^{\prime}$ and adjacency $A^{\prime}$, respectively. Subsequently, SEAL applies MPNN to this subgraph and use the sum of node representations within it, $\sum_{u}\text{MPNN}(u,A^{\prime},X^{\prime})$, as the target link representation. Other models employ a distinct approach to incorporate structural features. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023), for instance, directly apply MPNN to the original graph and concatenate structural features, such as the count of common neighbors, with the Hadamard product of target node MPNN representations, $\text{MPNN}(i,A,X)\odot\text{MPNN}(j,A,X)|\!|\text{structural features}$.
3.2 Structural Feature
Structural features for link prediction vary among models but are generally based on neighborhood overlap. Notable heuristics, such as Common Neighbor (CN), Resource Allocation (RA), and Adamic Adar (AA), use first-order common neighbors to compute scores for a target link $(i,j)$:
@@ 3.2 Structural Feature
3.3 Incompleteness of Graph
The primary aim of the link prediction task is to forecast unobserved edges, inherently making the input graph incomplete. Nevertheless, graph incompleteness can significantly impact structural features, such as common neighbors, and models based on them. This issue has drawn attention in some prior works. Yang et al. (2022) examined how unobserved links could distort evaluation scores, with a specific focus on metrics and benchmark design, while our research concentrates on model design. Dong et al. (2022) explored the consequences of the presence of target links, whereas our emphasis lies in understanding how incompleteness affects common neighbor-based features. Outside the domain of link prediction, Zhao et al. (2023) and Zhao et al. (2021) add edges predicted by GAE to the input graph of GNNs. However, their primary objective was node classification tasks, aiming to enhance edges between nodes of the same class while diminishing others. In contrast, our research addresses distribution shifts and information loss stemming from graph incompleteness, offering unique completion methods and insights tailored for link prediction.
@@ 5.1 Incompleteness Visualization
Given the pivotal role of common neighbor information in our NCN model and other link prediction models, we visualize the distribution of the number of common neighbors for training/test edges in both complete and incomplete graphs separately in Figure 4 (a)(c). To assess how incompleteness impacts model performance, we present the performance of CN model (as shown in Section 3.2) in four distinct scenarios in Figure 4 (b) (d). We observe the following effects of incompleteness:

### v4

@@ 1 Introduction
However, models combining SF and MPNN still have much room for improvement. They can generally be concluded into two architectures: SF-then-MPNN and SF-and-MPNN, as shown in Figure 2. For example, SEAL (Zhang & Chen, 2018) adds target-link-specific hand-crafted features to the node features of the input graphs of MPNN, whose output node representations are pooled to represent links. SEAL belongs to the SF-then-MPNN architecture, using SF to augment the input graph of MPNN. Though SF-then-MPNN models achieve provably high expressivity (Zhang et al., 2021), they require running MPNN on a different graph for each target link, leading to significant computational overhead. In contrast, Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) decouple the structural feature from MPNN, directly incorporating manually created pairwise features with individual node representations produced by MPNN as link representations, necessitating only a single run of MPNN on the original graph. These models fall under the SF-and-MPNN category, where structural features and MPNN are independent. Such methods have limited expressivity. For example, SEAL can capture the representations of common neighbor nodes, while BUDDY can only count the number of common neighbors.
To solve the drawback of two architectures above, we propose MPNN-then-SF architecture, which initially applies MPNN to the original graph and then uses structural features to guide the pooling of node representations. This approach offers strong expressivity and scalability: similar to SF-then-MPNN models, MPNN-then-SF can capture the node features of common neighbors, and similar to SF-and-MPNN models, it runs MPNN only once for all target links. We introduce the Neural Common Neighbor (NCN) as an instantiation of the MPNN-then-SF architecture. In experiments, NCN outperforms existing models in both scalability and performance.
@@
  • We introduce Neural Common Neighbor (NCN) with MPNN-then-SF architecture for link prediction, demonstrating superior performance and scalability compared to existing models.
@@ 2 Preliminary
We consider an undirected graph ${\mathcal{G}}\!=\!(V,E,A,X)$, where $V\!=\!\{1,2,\ldots,n\}$ is a set of $n$ nodes, $E\subseteq V\times V$ is the set of edges, $X\in\mathbb{R}^{n\times F}$ is a node feature matrix whose $v$-th row $X_{v}$ is the feature of node $v$, and a symmetric adjacency matrix $A\in\mathbb{R}^{n\times n}$ has $(u,v)$ element $A_{u}v=1$ if $(u,v)\in E$ and 00 otherwise. The degree of node $u$ is $d(u,A):=\sum_{v=1}^{n}A_{uv}$. Node $u$’s neighbors are nodes connected to $u$, $N(u,A):=\{v|v\!\in\!V,A_{uv}>0\}$. For simplicity, we use $N(u)$ to denote $N(u,A)$ when $A$ is fixed. Common neighbor means nodes connected to both $i$ and $j$: $N(i)\cap N(j)$.
@@
We define $A^{l}$ as a high-order adjacency matrix, where $A_{uv}^{l}$ represents the number of walks of length $l$ between nodes $u$ and $v$. $N(u,A^{l})=\{v|v\in V,A^{l}_{uv}>0\}$ denotes the set of nodes connected to $u$ by a walk of length $l$ in graph $A$. $N_{l}(u,A)$ denotes the set of nodes whose shortest path distance to $u$ in graph $A$ is $l$. Existing works define high-order neighbors as either $N(u,A^{l})$ or $N_{l}(u,A)$. More generally, the neighborhood of $u$ can be expressed as $N_{l_{1}}(u,A^{l_{2}})$, returning all nodes with a shortest path distance of $l_{1}$ to $u$ in the high-order graph $A^{l_{2}}$. We use $N_{l_{1}}^{l_{2}}(u)$ to denote $N_{l_{1}}(u,A^{l_{2}})$ when $A$ is fixed. Given a target link $(i,j)$, their general neighborhood overlap is given by $N_{l_{1}}^{l_{2}}(i)\cap N_{l_{1}^{\prime}}^{l_{2}^{\prime}}(j)$, and their neighborhood difference is given by $N_{l_{1}}^{l_{2}}(i)\!-\!N_{l_{1}^{\prime}}^{l_{2}^{\prime}}(j)$.
@@ 3 Related Work
Link Prediction Model
There are three primary categories of link prediction models. Node embeddings (Perozzi et al., 2014; Grover & Leskovec, 2016; Tang et al., 2015) map each node to an embedding vector and combine target nodes’ embeddings to predict link. Link prediction heuristics (Liben-Nowell & Kleinberg, 2003; Barabási & Albert, 1999; Zhou et al., 2009; Adamic & Adar, 2003) use hand-crafted structural features. GNNs utilize Graph Neural Networks. Among these GNNs, Graph Autoencoder (GAE) (Kipf & Welling, 2016) uses the inner product of two target nodes’ MPNN representations, $\langle\text{MPNN}(i,A,X),\text{MPNN}(j,A,X)\rangle$, as link $(i,j)$’s representation. It uses MPNN only and fails to capture pairwise relations between nodes. In contrast, various GNNs combining MPNN and structural features (Zhang & Chen, 2018; Yun et al., 2021; Chamberlain et al., 2023) have achieved state-of-the-art performance. SEAL (Zhang & Chen, 2018) is representative. For a target link $(i,j)$, it initially adds each node’s shortest path distance to $(i,j)$ to node feature $X$ and extracts a $k$-hop subgraph, generating augmented node feature $X^{\prime}$ and adjacency $A^{\prime}$. Then SEAL applies MPNN to the subgraph and pools node representations within it, $\sum_{u}\text{MPNN}(u,A^{\prime},X^{\prime})$, as the target link representation. Other models employ a distinct approach to incorporate structural features. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023), for instance, directly apply MPNN to the original graph and concatenate structural features, such as the count of common neighbors, with the Hadamard product of target node MPNN representations, $\text{MPNN}(i,A,X)\odot\text{MPNN}(j,A,X)|\!|\text{structural features}$.
Structural Feature
Structural features for link prediction vary but are generally based on neighborhood overlap. Notable heuristics, such as Common Neighbor (CN), Resource Allocation (RA), and Adamic Adar (AA), use first-order common neighbors to compute scores for a target link $(i,j)$:
@@
Incompleteness of Graph
Link prediction task is to forecast unobserved edges, inherently making the input graph incomplete. Nevertheless, graph incompleteness can significantly impact structural features, such as common neighbors, and models based on them. This issue has drawn attention in some prior works. Yang et al. (2022) examined how unobserved links could distort evaluation scores, with a specific focus on metrics and benchmark design, while our research concentrates on model design. Dong et al. (2022) explored the consequences of the presence of target links, whereas our emphasis lies in understanding how incompleteness affects common neighbor-based features. Outside the domain of link prediction, Zhao et al. (2023) and Zhao et al. (2021) add edges predicted by GAE to the input graph of GNNs. However, their primary objective was node classification tasks, aiming to enhance edges between nodes of the same class while diminishing others. In contrast, our research addresses distribution shifts and information loss stemming from graph incompleteness, offering unique completion methods and insights tailored for link prediction.
@@ 5.1 Incompleteness Visualization
Given the pivotal role of common neighbor information in our NCN model and other link prediction models, we visualize the distribution of the number of common neighbors for training/test edges in both complete and incomplete graphs separately in Figure 4 (a)(c). To assess how incompleteness impacts model performance, we present the performance of CN model (as shown in Section 3) in four distinct scenarios in Figure 4 (b) (d). We observe the following effects of incompleteness: